# DSCA: A Digital Subtraction Angiography Sequence Dataset and Spatio-Temporal Model for Cerebral Artery Segmentation

### v1

@@ Abstract
Cerebrovascular diseases (CVDs) remain a leading cause of global disability and mortality. Digital Subtraction Angiography (DSA) sequences, recognized as the golden standard for diagnosing CVDs, can clearly visualize the dynamic flow and reveal pathological conditions within the cerebrovasculature. Therefore, precise segmentation of cerebral arteries (CAs) and classification between their main trunks and branches are crucial for physicians to accurately quantify diseases. However, achieving accurate CA segmentation in DSA sequences remains a challenging task due to small vessels with low contrast, and ambiguity between vessels and residual skull structures. Moreover, the lack of publicly available datasets limits exploration in the field. In this paper, we introduce a DSA Sequence-based Cerebral Artery segmentation dataset (DSCA), the first publicly accessible dataset designed specifically for pixel-level semantic segmentation of CAs. Additionally, we propose DSANet, a spatio-temporal network for CA segmentation in DSA sequences. Unlike existing DSA segmentation methods that focus only on a single frame, the proposed DSANet introduces a separate temporal encoding branch to capture dynamic vessel details across multiple frames. To enhance small vessel segmentation and improve vessel connectivity, we design a novel TemporalFormer module to capture global context and correlations among sequential frames. Furthermore, we develop a Spatio-Temporal Fusion (STF) module to effectively integrate spatial and temporal features from the encoder. Extensive experiments demonstrate that DSANet outperforms other state-of-the-art methods in CA segmentation, achieving a Dice of 0.9033.
@@ I Introduction
Cerebrovascular diseases (CVDs) are significant causes of death, resulting in considerable suffering for patients and placing substantial financial pressure on families [1]. The majority of CVDs, such as stroke, moyamoya disease, and cerebral aneurysms, affect the geometrical and topological conditions of cerebral arteries (CAs). Therefore, analyzing CAs by segmenting and classifying them into bifurcated vessels (BV) and the main artery trunk (MTA) is essential for revealing vascular abnormalities associated with these diseases, leading to timely diagnosis and treatment planning [2].
In clinical practice, several imaging techniques such as computed tomography angiography (CTA), magnetic resonance angiography (MRA), and digital subtraction angiography (DSA) are commonly used for the diagnosis and treatment of CVDs. Among them, DSA offers dynamic imaging of cerebral vessels with high spatial and temporal resolution, providing accurate details of lesions [3]. It is thus commonly used as the golden standard for diagnosing CVDs by clinicians, in cases where CTA and MRA fail to provide accurate diagnosis. However, the examination using DSA heavily relies on visual inspection by radiologists, which can be time-consuming and labor-intensive. Furthermore, radiologists with limited experience may be susceptible to misdiagnosis and underdiagnosis. Therefore, automated segmentation of CAs in DSA sequences, with the capability of capturing precise cerebrovascular details, offers valuable features to expedite the diagnosis and treatment.
Figure 1: (a) is a DSA sequence, which includes 17 frames in the arterial phase. (b) is the MIP image of the arterial phase of the DSA sequence.
@@
Nevertheless, there is limited research on CA segmentation in DSA images. Zhang et al. [26] first proposed a straightforward U-shaped network for single-frame segmentation of DSA cerebral vessels. Meng et al. [7] introduced the MDCNN framework, leveraging CNN with multi-scale dense connections in conjunction with an entropy-sampled patch method for single-frame DSA segmentation. Xu et al. [27] proposed an edge regularization network (ERNet) for segmenting vessels in DSA images. ERNet uses erosion and dilation on the initial binary annotations to create pseudo-ground-truths for false negative and false positive cases. Subsequently, Vepa et al. [28] proposed a weakly supervised approach employing an active contour model to generate pseudo-labels for DSA vessel segmentation. However, despite these advancements, the segmentation performance of these methods still suffers from the low contrast of small vessels, and the interference from residual skull structures. Additionally, these methods focus only on pixel-by-pixel vessel extraction, while further classifying the vessel into bifurcated vessels (BV) and main artery trunk (MTA) is more clinically significant, e.g., for automatic detection and evaluation of stenosis, and automatic scoring for thrombolysis of cerebral infarction [29]. Moreover, regardless of whether fully supervised or weakly supervised methods are employed, existing approaches typically confine training on a single frame extracted from DSA sequences, leading to unstable microvascular segmentations.
@@
In clinical practice, a DSA sequence displays dynamic flows of vessels, where each frame captures only a fraction of the contrast agent, as depicted in Fig. 1-(a). Moreover, pathological changes such as arterial stenosis, collateralization, and aneurysms associated with various diseases exhibit significant variability across DSA sequences, appearing at uncertain frames within the sequence and manifesting diverse morphological characteristics. Consequently, relying solely on a single DSA frame for segmentation fails to cover the entire cerebrovasculature and these pathological features adequately. Unfortunately, a publicly available dataset that comprehensively labels entire dynamic blood flows is currently absent, primarily due to the labor-intensive nature of annotating such a sequence compared to labeling an individual frame. As illustrated in Fig. 1, the minimum intensity projection (MIP) image reveals the complete angiogram, incorporating information from all frames. These observations motivate us to consider training on entire DSA sequences and performing labeling on the MIP image. This approach not only preserves the complete cerebrovascular features but also alleviates the annotation workload simultaneously.
@@
Therefore, in this work, we establish a DSA Sequence dataset, comprising 224 DSA sequences of several common CVDs. To address the segmentation challenges, we propose a spatio-temporal network (DSANet) for CA segmentation. It takes a DSA sequence and its MIP image as input to produce a complete 2D segmentation. Unlike 2D segmentation methods that only use a spatial encoding branch (SEB), our network includes an independent temporal encoding branch (TEB) to capture crucial temporal information. Following the TEB, we treat each frame in the sequence as a token, leveraging TemporalFormer blocks to capture global flow information from the DSA sequence. Subsequently, the spatial and temporal features from the two encoding branches are effectively fused through the spatio-temporal fusion (STF) module. In summary, our main contributions can be outlined as follows:
@@
$\bullet$ We establish the first publicly accessible DSA-Sequence based Cerebral Artery segmentation dataset (DSCA), consisting of pixel-wise cerebrovascular annotations across 224 DSA sequences from several common CVDs. To promote further developments in this field, the DSCA dataset, the code, and baseline models are made available at this website.
@@
$\bullet$ We propose a novel spatio-temporal strategy aimed at enhancing cerebral artery segmentation within DSA sequences. We design two encoding branches, i.e., the TEB and SEB, which respectively employ the sequence images and MIP images as input, to leverage both spatial and temporal information for precise segmentation of the entire cerebrovasculature.
@@
$\bullet$ We introduce a new STF module to integrate spatial and temporal features from the encoders. This module effectively addresses ambiguity among BV, MTA, and residual skulls by leveraging the dynamic changes of vessels and the structural invariance of the skull.
@@ II-A Dataset Description
The DSCA dataset consists of 58 patients, in total 224 DSA sequences including 1792 images from the left and right hemispheres. Among them, there are 28 male and 30 female patients, with ages ranging from 9 to 81 years and an average age of 49. Notably, the DSA sequences include three different arteries: the internal carotid artery (ICA), external carotid artery (ECA), and vertebral artery (VA). Specifically, there are 126 sequences for ICA, 55 for ECA, and 43 for VA. These sequences were captured by multiple imaging devices including AXION-Artis-160145 and AXION-Artis-160480 (Siemens, Germany), Azurion-559, Azurion-703844, AlluraXper-722012-2542 and AlluraXper-722038-129 (Philips, Netherlands), and Bransist Safire VC17 (Shimadzu, Japan), with sampling rates ranging from 4 to 7 frames per second. Moreover, these sequences exhibit different resolutions (from $512\times 472$ to $1432\times 1432$). Each DSA sequence, captured in coronal or sagittal views, was stored in a DICOM file, with 108 DSA sequences in the coronal view and 116 in the sagittal view. All sequences retained arterial phase frames while discarding non-contrast, capillary phase, and venous phase frames [29], guided by the neurosurgeon’s expertise. Consequently, DSA sequences in the DSCA dataset consist of 5 to 22 frames. To align with network input, we resample all DSA sequences to 8 frames based on the distribution of extracted arterial phase frames. Detailed information about the dataset is provided in the GitHub.
TABLE I: Dataset comparison
@@
Figure 2: The annotating examples of the DSCA dataset. The first to third rows are ICA, ECA, and VA. The 1st and 2nd column are coronal view, and the 3rd and 4th column are sagittal view.
@@ II-B Manual Annotation Protocol
As mentioned above, manually annotating the complete DSA sequence is a time-consuming and labor-intensive task. In contrast, annotating the corresponding MIP image can save considerable time while retaining the entire cerebrovascular information. Furthermore, distinguishing between BV and MTA holds greater clinical significance. Thus, we asked skilled clinicians to manually annotate the MIP images, including specific labeling for both BV and MTA. These annotations served as the gold standard for the DSCA dataset. The data preprocessing and annotation procedures are as follows: (1) Selection of arterial phase frames from the DSA sequences. (2) Registration of the extracted arterial phase sequence frames using Elastix [31], followed by resampling to 8 frames. (3) Projection of the resampled frames to generate MIP images. (4) Annotation of the MIP images utilizing ITK-SNAP [32].
The annotation process of the DSCA dataset was performed by five skilled clinicians who received complete training. Disputed annotations were referred to a senior neurosurgeon for confirmation. Subsequently, a second senior neurosurgeon conducted a final review and revision of all annotations. The annotated results of the DSCA dataset are presented in Fig. 2, wherein the ICA and VA sequences were categorized into background, BV, and MTA. The ECA sequences, lacking a distinct main trunk, were classified into background and BV. To mitigate potential risks of data leakage, the DSA sequences were randomly divided based on patient IDs into a training set comprising $80\%$ of the data and a test set comprising the remaining $20\%$. Consequently, the training set comprises 180 sequences, while the test set contains 44 sequences. Table I provides a detailed comparison of the DSCA dataset and other DSA datasets. Notably, the DSCA dataset offers several distinct advantages, including a large number of samples, diverse image resolutions, and most importantly, public accessibility.
Figure 3: Architecture of our proposed DSANet. It is primarily composed of two encoding branches, one decoding branch, and two modules: TemporalFormer (TF), Spatio-Temporal Fusion (STF).
@@ III-A Encoding Branches
As shown in Fig. 3, there are two encoding branches: TEB at the top and SEB at the bottom. Although structurally similar, these branches function independently without weight sharing. Each encoding branch adopts a five-layer structure, comprising a single convolutional layer followed by four downsampling layers. The downsampling layer comprises a Maxpool block and a convolutional layer with double $3\times 3$ convolutional blocks, double GroupNorm blocks, and double Gelu blocks. The primary difference between the two encoding branches lies in their input.
@@ III-A1 Spatial encoding branch
In the SEB, the input is a 2D MIP image $M\in\mathbb{R}^{B\times C\times H\times W}$, where $B$, $C$, $H$, and $W$ represent the batch size, number of channels, height, and width of $M$, respectively. This input setting is consistent with other 2D segmentation tasks in our subsequent comparisons.
@@ III-A2 Temporal encoding branch
In the TEB, the input is organized in 2D sequential frames denoted as $S\in\mathbb{R}^{B\times T\times C\times H\times W}$, where $T$ represents the number of frames. The sequence encoding process is optimized for training convenience by merging $B$ and $T$ into $N$, where $N=B\times T$. Following the first convolutional layer, the value of $C$ is doubled. Afterward, as the input sequence passes through the downsampling layer, both $H$ and $W$ are halved, and $C$ is doubled. Each of the first four layers of the encoding branches includes an additional output serving as a skip connection. In the TEB, the skip connection needs to be downsampled in the $T$ dimension and then concatenated with the skip connection in the SEB to form the final skip connection, and is subsequently fed into the corresponding layer in the decoder.

### v2

@@ Abstract
Cerebrovascular diseases (CVDs) remain a leading cause of global disability and mortality. Digital Subtraction Angiography (DSA) sequences, recognized as the gold standard for diagnosing CVDs, can clearly visualize the dynamic flow and reveal pathological conditions within the cerebrovasculature. Therefore, precise segmentation of cerebral arteries (CAs) and classification between their main trunks and branches are crucial for physicians to accurately quantify diseases. However, achieving accurate CA segmentation in DSA sequences remains a challenging task due to small vessels with low contrast, and ambiguity between vessels and residual skull structures. Moreover, the lack of publicly available datasets limits exploration in the field. In this paper, we introduce a DSA Sequence-based Cerebral Artery segmentation dataset (DSCA), the publicly accessible dataset designed specifically for pixel-level semantic segmentation of CAs. Additionally, we propose DSANet, a spatio-temporal network for CA segmentation in DSA sequences. Unlike existing DSA segmentation methods that focus only on a single frame, the proposed DSANet introduces a separate temporal encoding branch to capture dynamic vessel details across multiple frames. To enhance small vessel segmentation and improve vessel connectivity, we design a novel TemporalFormer module to capture global context and correlations among sequential frames. Furthermore, we develop a Spatio-Temporal Fusion (STF) module to effectively integrate spatial and temporal features from the encoder. Extensive experiments demonstrate that DSANet outperforms other state-of-the-art methods in CA segmentation, achieving a Dice of 0.9033.
@@ I Introduction
Cerebrovascular diseases (CVDs) are significant causes of death, resulting in considerable suffering for patients and placing substantial financial pressure on families [1]. The majority of CVDs, such as stroke, moyamoya disease, and cerebral aneurysms, affect the geometrical and topological conditions of cerebral arteries (CAs). Therefore, segmenting CAs is essential for revealing vascular abnormalities associated with these diseases, leading to timely diagnosis and treatment planning [2].
In clinical practice, several imaging techniques such as computed tomography angiography (CTA), magnetic resonance angiography (MRA), and digital subtraction angiography (DSA) are commonly used for the diagnosis and treatment of CVDs. Among them, DSA offers dynamic imaging of cerebral vessels with high spatial and temporal resolution, providing accurate details of lesions [3]. It is thus commonly used as the gold standard for diagnosing CVDs by clinicians, in cases where CTA and MRA fail to provide accurate diagnosis. However, the examination using DSA heavily relies on visual inspection by radiologists, which can be time-consuming and labor-intensive. Furthermore, radiologists with limited experience may be susceptible to misdiagnosis and underdiagnosis. Therefore, automated segmentation of CAs in DSA sequences, with the capability of capturing precise cerebrovascular details, offers valuable features to expedite the diagnosis and treatment.
Figure 1: (a) is a DSA sequence, which includes 17 frames in the arterial phase. (b) is the MinIP image of the arterial phase of the DSA sequence.
@@
Moccia et al. [26] and Mookiah et al. [27] summarize numerous vascular segmentation methods and datasets, which have significantly advanced the field of vessel segmentation. However, the majority of these methods are applied to retinal vessel segmentation. Compared to other imaging techniques, such as color fundus photography, acquiring DSA images is more challenging and involves higher risks. This has resulted in limited research on CA segmentation in DSA images. In DSA cerebrovascular segmentation, Zhang et al. [28] first proposed a straightforward U-shaped network for single-frame segmentation of DSA cerebral vessels. Patel et al. [29] compared the performance of UNet [8] and DeepMedic [30] in DSA cerebral vessel segmentation, demonstrating the effectiveness of DeepMedic’s use of multi-resolution input. Meng et al. [7] introduced the MDCNN framework, leveraging CNN with multi-scale dense connections in conjunction with an entropy-sampled patch method for single-frame DSA segmentation. Xu et al. [31] proposed an edge regularization network (ERNet) for segmenting vessels in DSA images. ERNet uses erosion and dilation on the initial binary annotations to create pseudo-ground-truths for false negative and false positive cases. Subsequently, Vepa et al. [32] proposed a weakly supervised approach employing an active contour model to generate pseudo-labels for DSA vessel segmentation. However, despite these advancements, the segmentation performance of these methods still suffers from the low contrast of small vessels, and the interference from residual skull structures. Additionally, these methods focus only on pixel-by-pixel vessel extraction, while further classifying the vessel into bifurcated vessels (BV) and main artery trunk (MAT) is more clinically significant, e.g., for automatic detection and evaluation of stenosis, and automatic cerebrovasculature labeling [33]. Moreover, regardless of whether fully supervised or weakly supervised methods are employed, existing approaches typically confine training on a single frame extracted from DSA sequences, leading to unstable microvascular segmentations.
@@
In clinical practice, a DSA sequence displays dynamic flows of vessels, where each frame captures only a fraction of the contrast agent, as depicted in Fig. 1-(a). Moreover, pathological changes such as arterial stenosis, collateralization, and aneurysms associated with various diseases exhibit significant variability across DSA sequences, appearing at uncertain frames within the sequence and manifesting diverse morphological characteristics. Consequently, relying solely on a single DSA frame for segmentation fails to cover the entire cerebrovasculature and these pathological features adequately.
As illustrated in Fig. 1, the minimum intensity projection (MinIP) image reveals the complete angiogram, incorporating information from all frames. These observations motivate us to consider training on entire DSA sequences and performing labeling on the MinIP image. This approach not only preserves the complete cerebrovascular features but also alleviates the annotation workload simultaneously. Moreover, our previous work [34] has demonstrated that temporal information can enhance cerebrovascular segmentation performance. Recently, Liu et al. [35] publicly released a DSA sequence dataset, DIAS. They established benchmarks for this dataset across three levels: fully supervised, weakly supervised, and semi-supervised, significantly advancing DSA research. The DIAS dataset contains 60 labeled samples and 60 unlabeled samples, and it lacks detailed descriptions of the pathological information.
@@
Therefore, in this work, we establish a DSA Sequence dataset, comprising 224 DSA sequences of several common CVDs. To address the segmentation challenges, we propose a spatio-temporal network (DSANet) for CA segmentation. It takes a DSA sequence and its MinIP image as input to produce a complete 2D segmentation. Unlike 2D segmentation methods that only use a spatial encoding branch (SEB), our network includes an independent temporal encoding branch (TEB) to capture crucial temporal information. Following the TEB, we treat each frame in the sequence as a token, leveraging TemporalFormer blocks to capture global flow information from the DSA sequence. Subsequently, the spatial and temporal features from the two encoding branches are effectively fused through the spatio-temporal fusion (STF) module. In summary, our main contributions can be outlined as follows:
@@
$\bullet$ We establish the new publicly accessible DSA-Sequence based Cerebral Artery segmentation dataset (DSCA), consisting of pixel-wise cerebrovascular annotations across 224 DSA sequences from several common CVDs. To promote further developments in this field, the DSCA dataset, the code, and baseline models are made available at this website [^1] https://github.com/jiongzhang-john/DSCA.
@@
$\bullet$ We propose a novel spatio-temporal strategy aimed at enhancing cerebral artery segmentation within DSA sequences. We design two encoding branches, i.e., the TEB and SEB, which respectively employ the sequence images and MinIP images as input, to leverage both spatial and temporal information for precise segmentation of the entire cerebrovasculature.
@@
$\bullet$ We introduce a new STF module to integrate spatial and temporal features from the encoders. This module effectively addresses ambiguity among BV, MAT, and residual skulls by leveraging the dynamic changes of vessels and the structural invariance of the skull.
TABLE I: Dataset description. Internal Carotid Artery (ICA), External Carotid Artery (ECA), Vertebral Artery (VA), Antero-Posterior (AP), Lateral (LA).
Artery Total View Devices Diseases
ICA 126 AP: 62 Siemens:36 Normal: 28
Philips: 84 Stenosis: 23
LA: 64 Shimadzu: 6 Moyamoya: 48
Aneurysm: 27
ECA AP: 26 Siemens: 30 Normal: 55
55 Philips: 22
LA: 29 Shimadzu: 3
VA AP: 17 Siemens: 17 Normal: 37
43 Philips: 23 Stenosis: 2
LA: 26 Shimadzu: 3 Moyamoya: 4
@@ II-A Dataset Description
The DSCA dataset consists of 58 patients, in total 224 DSA sequences including 1792 images from the left and right hemispheres. Among them, there are 28 male and 30 female patients, with ages ranging from 9 to 81 years and an average age of 49. Notably, the DSA sequences include three different arteries: the internal carotid artery (ICA), external carotid artery (ECA), and vertebral artery (VA). Specifically, there are 126 sequences for ICA, 55 for ECA, and 43 for VA. These sequences were captured by multiple imaging devices including AXION-Artis-160145 and AXION-Artis-160480 (Siemens, Germany), Azurion-559, Azurion-703844, AlluraXper-722012-2542 and AlluraXper-722038-129 (Philips, Netherlands), and Bransist Safire VC17 (Shimadzu, Japan), with sampling rates ranging from 4 to 7 frames per second, which is slightly different from what is reported in [36]. The disease and manufacturer information of each sequence are shown in Table I. Moreover, these sequences exhibit different resolutions (from $512\times 472$ to $1432\times 1432$). Each DSA sequence, captured in antero-posterior or lateral views, was stored in a DICOM file, with 108 DSA sequences in the antero-posterior view and 116 in the lateral view. All sequences retained arterial phase frames while discarding non-contrast, capillary phase, and venous phase frames [37], guided by the neurosurgeon’s expertise. Therefore, DSA sequences in DSCA dataset consist of 5 to 22 frames. To align with network input, we resample all DSA sequences to 8 frames based on the distribution of extracted arterial phase frames. Detailed information about the dataset is provided in the GitHubfootnotemark: .
TABLE II: Dataset comparison
@@
Patel et al. [29] 100 - Single Frame No
DIAS [35] 120 800 $\times$ 800 Sequence Yes
@@
Figure 2: The annotating examples of the DSCA dataset. The first to third rows are ICA, ECA, and VA. The 1st and 2nd columns are antero-posterior view, and the 3rd and 4th columns are lateral view.
@@ II-B Manual Annotation Protocol
As mentioned above, manually annotating the complete DSA sequence is a time-consuming and labor-intensive task. In contrast, annotating the corresponding MinIP image can save considerable time while retaining the entire cerebrovascular information. Furthermore, distinguishing between BV and MAT holds greater clinical significance. Thus, we asked skilled clinicians to manually annotate the MinIP images, including specific labeling for both BV and MAT. These annotations served as the gold standard for the DSCA dataset. The data preprocessing and annotation procedures are as follows: (1) Selection of arterial phase frames from the DSA sequences. (2) Registration of the extracted arterial phase sequence frames using Elastix [39], followed by resampling to 8 frames. (3) Projection of the resampled frames to generate MinIP images. (4) Annotation of the MinIP images utilizing ITK-SNAP [40].
The annotation process of the DSCA dataset was performed by five skilled clinicians with over 5 years of experience who received complete training. They were evenly assigned to annotate the 224 MinIP images, each being annotated only once. Disputed annotations were referred to a senior neurosurgeon for confirmation. Subsequently, a second senior neurosurgeon conducted a final review. If there were different opinions, the two neurosurgeons worked together to reach a consensus. The annotated results of the DSCA dataset are presented in Fig. 2, wherein the ICA and VA sequences were categorized into background, BV, and MAT. The ECA sequences, lacking a distinct main trunk, were classified into background and BV. The criteria for dividing MAT and BV are based on the last segment of the ICA and VA. The ICA is labeled up to the C7 segment, and the VA up to the V4 segment, merging with the basilar artery. To mitigate potential risks of data leakage, the DSA sequences were randomly divided based on patient IDs into a training set comprising $80\%$ of the data and a test set comprising the remaining $20\%$. This ensures that the same patient cannot appear in both the training and testing sets simultaneously. Consequently, the training set comprises 180 sequences, while the test set contains 44 sequences.
Additionally, Table II provides a detailed comparison of the DSCA dataset and other DSA datasets. Notably, the DSCA dataset offers several distinct advantages, including a large number of samples, diverse image resolutions, and most importantly, public accessibility. Even compared to the recently released DIAS dataset, which is also a DSA sequence dataset, DSCA shows clear advantages over DIAS in terms of sample size and diversity in image resolution.
Figure 3: Architecture of our proposed DSANet. It is primarily composed of two encoding branches, one decoding branch, and two modules: TemporalFormer (TF), Spatio-Temporal Fusion (STF). The details of the dimension transformation process of TF are shown in Fig. 4
@@ III-A Encoding Branches
As shown in Fig. 3, there are two encoding branches: TEB and SEB. Although structurally similar, these branches function independently without weight sharing. Each branch adopts a five-layer structure, comprising a convolutional layer with double $3\times 3$ convolutional blocks, double GroupNorm blocks, and double Gelu blocks, followed by four downsampling layers. The downsampling layer comprises a Maxpool block and a convolutional layer. The primary difference between the two encoding branches lies in their input.
@@ III-A1 Spatial encoding branch
In the SEB, the input is a 2D MinIP image $M\in\mathbb{R}^{B\times C\times H\times W}$, where $B$, $C$, $H$, and $W$ represent the batch size, number of channels, height, and width of $M$. And $C$ is equal to 1. This input setting is consistent with other 2D segmentation tasks in our subsequent comparisons.
@@ III-A2 Temporal encoding branch
In the TEB, the input is organized in 2D sequential frames denoted as $S\in\mathbb{R}^{B\times T\times C\times H\times W}$, where $T$ represents the number of frames. The sequence encoding process is optimized for training convenience by merging $B$ and $T$. Following the first convolutional layer, the value of $C$ is doubled. Afterward, as the input sequence passes through the downsampling layer, both $H$ and $W$ are halved, and $C$ is doubled. Each of the first four layers of the encoding branches includes an additional output serving as a skip connection. In the TEB, the skip connection needs to be downsampled in the $T$ dimension and then concatenated with the skip connection in the SEB to form the final skip connection, and is subsequently fed into the corresponding layer in the decoder.